OpenStudy (karatechopper):

this is for ishaan:3+2=

OpenStudy (osanseviero):

it is 10 in base 5

OpenStudy (karatechopper):

i said for ishaan!!

OpenStudy (anonymous):

\[\nu \pi \kappa \xi \pi \pi \pi \xi \xi \xi \xi \xi \xi \xi \epsilon \xi \xi \xi \xi \xi ^{FIVE}\rho \xi \iota \pi \pi \kappa \rho \lambda \lambda\]

OpenStudy (karatechopper):

not u!!

OpenStudy (anonymous):

ishaan is asleep.

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

MAYBE KILLED...

OpenStudy (karatechopper):

lol noooo!!

OpenStudy (anonymous):

lol maybe.

OpenStudy (karatechopper):

lol no

OpenStudy (anonymous):

*hides blood-stained knife*

OpenStudy (anonymous):

AHHHHHH!!!!!

OpenStudy (karatechopper):

not nice...

OpenStudy (karatechopper):

murder!!!

OpenStudy (anonymous):

GOODMAN is next. lawl

OpenStudy (anonymous):

poor ishaan...he was my admission officer for EFC

OpenStudy (anonymous):

NOOOO...IMA GOODMAN DUNT KILL ME!!!

OpenStudy (karatechopper):

efc?

OpenStudy (anonymous):

EPIC FAIL CLUB ...comon dunt tell me u forgot

OpenStudy (karatechopper):

i never knew about it

OpenStudy (anonymous):

wat are u running for? watt? we were talking bout it for ages

OpenStudy (karatechopper):

o!!! i kno now!! what am i?? the..

OpenStudy (anonymous):

run for watever u want..so far, prez, vice prez, janitor, treasurer, and admission police is taken

OpenStudy (anonymous):

so wat are u gonna run for?

OpenStudy (karatechopper):

aww darn it i was the treasurer!!! i already have a post!!

OpenStudy (anonymous):

sum guy took it...u could be secretary or senator...

OpenStudy (anonymous):

or jus make a funnay one...

OpenStudy (karatechopper):

im the senator i say yes to votes

OpenStudy (anonymous):

okay, ur elected as senator for EFC!!! jus tell ishaan so he could put u in the book of members....

OpenStudy (karatechopper):

hes asleep

OpenStudy (anonymous):

well...wn he comes bak!! LOL...

OpenStudy (karatechopper):

i hope!! unless insanity guy killed him:)

OpenStudy (anonymous):

nooo dunt say tat!!!!

OpenStudy (anonymous):

everyone is dying its just a matter of when.

OpenStudy (anonymous):

AHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH

OpenStudy (karatechopper):

lol

OpenStudy (anonymous):

WHY? WAT DID ANYBODY DO TO U?

OpenStudy (anonymous):

I'm the serial killer on OpenStudy.

OpenStudy (karatechopper):

better put tht as an abuse then!!

OpenStudy (anonymous):

I become better at math when I take their souls.

OpenStudy (anonymous):

u gotta have a reason, is it depression? i noe a good therapist....

OpenStudy (anonymous):

rofl

OpenStudy (anonymous):

we know..depression hurts, but u dunt have to take the lives of innocent ppll!!!

OpenStudy (anonymous):

32

OpenStudy (karatechopper):

i belive its 5... go back to KG

OpenStudy (anonymous):

32

OpenStudy (karatechopper):

I like how ishaan just came in, did his buisness of answering the question, and walked out.

OpenStudy (anonymous):

Wow.